9. e in pledges of David is answering the telephone during a radio broadcast fund drive. People phon $5 or more per month. Experience has shown that 45% of the callers pledge more th month. an $5 per (a) What is the probability that on a given day the first caller to pledge over $5 per month is the fourth call of the day? (b) What is the probability that the first caller to pledge over $5 per month will be among the first three calls?
Answer #1

a)
p = 0.45

Required proability = (1-0.45)^3 * 0.45 = 0.0749

b)
Required probability = 3C1 * 0.45 * 0.55^2 = 0.4084